A proton is traveling horizontally to the right at 4.80×106 m/s

Find (a)the magnitude and (b) direction of the weakest electric field that can bring the proton uniformly to rest over a distance of 3.10 cm .

(c)How much time does it take the proton to stop after entering the field?

What minimum field ((d)magnitude and (e)direction) would be needed to stop an electron under the conditions of part (a)?
Answer #1

a)

using 3rd equation of motion

v^2 =u^2 - 2 a d

0 = ( 4.8*10^6)^2 - 2* a* 0.031

a = 3.716*10^14 m/s^2

now applying equilibrium of forces

qE = ma ... {i)

E = 1.67*10^-27* 3.716*10^14 / ( 1.6*10^-19)

E = 3.88*10^6 N/C

directed opposite to the direction of motion of proton, i.e. to the left.

==========

t = v/a = 4.8*10^6/ (3.716*10^14)

t = 1.292*10^-8 seconds

=========

for electron

Considering eqn .. (i)

E = ma / q

E = 9.1*10^-31* 3.716*10^14 / ( 1.6*10^-19)

E = 2113.475 N/C

directed to the right.

===========

do comment in case any doubt, will reply for sure.. Goodluck